OpenStudy (anonymous):

If f(x) = 2x + 6, find f(7).

hero (hero):

f(x) = 2x + 6 f(7) = 2(7) + 6 f(7) = 14 + 6 f(7) = 20

OpenStudy (anonymous):

gee if I split each of the seven logs...